OpenStudy (anonymous):

If a triangle has an area of 56sq. inches, with base 9 inches longer than its altitude, find the length of its base.

OpenStudy (a_clan):

0.5 * x * (x-9) = 56

OpenStudy (anonymous):

then ? whats next ?

OpenStudy (saifoo.khan):

solve for x for find it's base

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

saifoo.khan : can you complete the sol'n . thanks a lot :))

OpenStudy (saifoo.khan):

then you will get base.

OpenStudy (saifoo.khan):

Alright, sure.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

0.5 * x * (x-9) = 56 x(x-9) =56/0.5 x^2 -9x = 112 x^2 -9x -112 = 0 Solve thequadratic equation, the solution which u will get will be: x = 16, x = -7

OpenStudy (saifoo.khan):

did you get the answer?

OpenStudy (anonymous):

i think 25 not 15 ?

OpenStudy (saifoo.khan):

Oh, my bad. :/

OpenStudy (saifoo.khan):

-7 will be neglected, so the altitude is 16. And the base is (16+9)=25units.

OpenStudy (saifoo.khan):

i hate such typos! >_<

OpenStudy (anonymous):

its okay .. thanks a lot .. can answer this one .. pls ??

OpenStudy (saifoo.khan):

this one?

OpenStudy (anonymous):

a rectangular cardboard is to be made into an open box by cutting a 2-cm sq. piece for each corner and turning up the sides. If the perimeter of the cardboard is 60 cm and the volume of the resulting box is 224 cm^3, what are the dimensions of the cardboard?

OpenStudy (anonymous):

pls ?

OpenStudy (saifoo.khan):

ok, i have to get going to grab something to drink, i will br back in 2mins! :)